4. Divide the number 32 in two parts so that one part is 3 times larger than the other?
Answer by ewatrrr(24785)


Hi
one part is 3 times larger than the other
Let x and 3x represent those numbers
x + 3x = 32
4x = 32
x = 8, smaller of those two numbers. larger = 24 (3*8)
8 + 24 = 32